OpenStudy (anonymous):

x is an integer and 9 < x2 < 99. What is the maximum value of x minus the minimum value of x?

OpenStudy (jango_in_dtown):

\[9<x^2<99 \]

OpenStudy (jango_in_dtown):

implies \[-3<x<\sqrt{99}\]

OpenStudy (jango_in_dtown):

so max -min = \[\sqrt{99}-(-3)=\sqrt{99}+3=3(1+\sqrt{11})\]